OpenStudy (monteskellie):

What is the value of y in the solution to the following system of equations? 5x − 3y = −3 2x − 6y = −6

OpenStudy (monteskellie):

o, -1, 1, 2 are my choices for an answer.

OpenStudy (monteskellie):

I believe it is 1

OpenStudy (wolf1728):

a) 5x − 3y = −3 b) 2x − 6y = −6 Multiply equation b by -2.5 a) 5x − 3y = −3 b) -5x +15y = 15 Add the equations: 12 y = -12

OpenStudy (silverfang492):

First, multiply the first equation by 2, so you have 10x-6y=-6. Then, you multiply the bottom equation by 5 which gives you 10x-30y=-30. After that, you subtract the bottom equation from the top equation which gives you 24y=24. Finally, you divide both sides by 24 which lieaves you with y=1.

OpenStudy (wolf1728):

I made a mistake in my calculations. Adding the equations produces 12 y = 12 and y = 1 silverfang492 is correct

OpenStudy (silverfang492):

yeah you accidentally said -12 instead of 12